A train leaves the station at 12:00 travling 90 mph and travels 450 miles to its destination. Another train leaves at 12:15 traveling 100 mph from the 1st train's destination. What time will they possibly pass each other?
Answer by stanbon(75887)

A train leaves the station at 12:00 travling 90 mph and travels 450 miles to its destination. Another train leaves at 12:15 traveling 100 mph from the 1st train's destination. What time will they possibly pass each other?
------------------------------
1st train DATA:
rate = 90 mph ; distance = x miles ; time = x/90 hrs.
----------------------
2nd train DATA:
rate = 100 mph ; distance = 450-x miles ; time = (450-x)/100 hrs
--------------------------
Equation:
1st train time - 2nd train time = 1/4 hr.
------
x/90 - (450-x)/100 = 1/4
------
Multiply thru by 900 to get:
10x - 9(450-x) = 225
10x - 9*450 + 9x = 225
19x = 19*225
x = 225 miles (1st train distance)
1st train time = 225/90 = 2.5 hrs
----
Clock time when they pass one-another: 2:30 PM
==============================
Cheers,
Stan H.